what is the slope and y intercept of 6-1/3=7x?
Answer by Alan3354(69443)

what is the slope and y intercept of 6-1/3=7x?
---------
6-1/3=7x
7x = 17/3
x = 17/21
----------
This is a "vertical" line thru (17/21,0). Its slope is infinite, or undefined. There is no y-intercept, it's parallel to the y-axis.